OpenStudy (anonymous):

She completed the first session in 56 minutes, the second in 42 minutes and the third in 1 hour and 16 minutes. How many hours did it take her to complete all three sessions?

Parth (parthkohli):

56 + 42 + 76 = 100 + 76 = 176 minutes. 176 minutes = 2 hours 56 mins.